OpenStudy (anonymous):

Which mixed number is equivalent to thirteen-fifths? one and two-fifths one and three-fifths two and three-fifths three and one-fifth

jimthompson5910 (jim_thompson5910):

Hint: 13/5 = (10+3)/5 = 10/5 + 3/5